-How many O2 molecules are needed to react with 2.80 g of S?
-What is the theoretical yield of SO3 produced by 2.80 g of S?
1
Expert's answer
2020-10-30T06:46:08-0400

Reaction of S with O2 is :

 2S(g) + 3O2(g) –––> 2SO3(g)

  

According to above reaction

2 mole of S needs 3 moles of O2


Molar mass of S = 32 g/mol

So, mass of 2 mole of S = 32 g/mol × 2 mole

                      = 64 g 

So, 

64 g of S needs 3 moles of O2

1 g of S needs (3/64) mole of O2

2.80 g of S needs (3/64)×2.80 mole of O2

               = 0.1313 mole O2



Part 1            

In 1 mole of any substance number of molicule = 6.023×1023 molicules


So, 0.1313 mole O2

    = (0.1313×6.023×10^23) molicules

    = 7.908×1022 molicules


Hence,

7.908×1022 molicules O2 molecules are needed to react with 2.80 g of S



Part 2

Theoriticly 

2.80 g of S needs 0.1313 mole O2


Molar mass of O2 = 32 g/mol

Mass of 0.1313 mole O2 = 32×0.1313 g

                      = 4.202 g 

Hence,

Theoriticly 2.80 g of S needs 4.202 g O2


So, theoriticl yield = 4.202 g